# The Lunar Gravitational-wave Antenna: Mission Studies and Science Case

### v1

@@ 3.2.1 Astrophysical explosions and matter effects on waveforms
In the last two decades, significant efforts have been made in investigating the GW emission from CCSNe (see [304, 305, 306] for recent reviews). Asymmetries of the explosion are required to produce GWs. This happens during three distinct phases: a) immediately at bounce (time-dependent rotational flattening and post-shock convection, lasting tens of ms); b) a low-state in which the neutrino-driven turbulence grows (lasting 50–200 ms); and through c) Proto-NS modal oscillations [307, 308, 309, 310, 311, 312, 313, 314, 304]. These phases, characterised by GW emission in the $\sim$100–1000 Hz frequency range, are accompanied by an additional low-frequency ($\lesssim$50 Hz) signal, the so-called GW memory [315] (see also Sec. 3.2.2). This is generated by asymmetric neutrino emission and aspherical, large-scale matter ejecta motions [316, 317, 318, 319, 320, 321, 322]. The neutrino memory signal shows a more secular time-evolution than the matter component, which is due to the cumulative time-integral of the anisotropy-weighted neutrino luminosity [323]. Sustained turbulent accretion in more massive progenitors results in higher neutrino luminosities and generally more anisotropic explosions. As with the matter component, the neutrino component is more pronounced for delayed explosions of models with higher compactness [324].
@@
Large-scale SN ejecta asymmetries (where jet-like structures may emerge) are expected (see e.g. [325, 326, 327, 328, 329]) and also supported observationally by spectropolarimetry [330, 331, 332, 333, 334, 335, 336], but their GW emissions are associated with frequencies that are beyond the LGWA observation band (100–1000 Hz) [337, 338, 339, 340] (see also Fig. 27).
@@
Figure 27: Gravitational-wave spectra $2\sqrt{f}|\tilde{h}(f)|$ for CCSNe with different stellar masses as a function of time projected against strain noise curves $\sqrt{S_{n}}$ for various GW detectors (adapted from [306]). The simulations, extending out to $\sim$6 seconds after the bounce, are unable to constrain the spectrogram in the region 0.1–1.0 Hz.
@@
While most of the published simulations do not cover the decihertz band (see for instance [306] and Fig. 27) the linear and nonlinear memory effects can lead to detectable GW signals below 1 Hz [341, 321]. In this way, even with LGWA, one would be able to study the GW emission generated by the neutrino processes [342, 319, 306]. As discussed above, the range $>$25 Hz has signals from the explosion asymmetries, and the range $<$10 Hz carries signals from the anisotropy of the neutrino emissions. In principle, these are distinguishable (though they no doubt overlap) and the polarization and signal together say something about the various asymmetries of the event and the orientation of the sources with respect to the line of sight. In this respect, it is worth noticing that a simultaneous detection of the low-frequency memory signal by LGWA and the high-frequency burst by terrestrial detectors would provide a more complete understanding of CCSNe.
Although exploration of the lower-frequency CCSN signals and their detectability have only started in the recent years, the progress made so far in extending the simulations to several seconds after the bounce, and the immediate future prospects of reaching out to tens of seconds, carry the promise of producing consistent results at low frequencies.
@@
The acceleration of a relativistic jet is another source of gravitational radiation [343, 344, 345]. The resulting GW signal is a memory type one and its detailed features depend on the acceleration (and possible deceleration) of the jet, on its overall duration and orientation relative to the observer [346, 347, 348]. The overall amplitude of the GW signal is, as expected on dimensional grounds, of order $GE_{\rm jet}/c^{4}d$, where $E_{\rm jet}$ is the energy of the jet and $d$ the distance of the source. The frequency is the smaller of the inverse acceleration time and the jet duration. Natural candidates for jet-GWs are long GRB jets whose typical energy is of order $10^{51}$ erg and typical frequency of order tens of deciHz. Short GRB jets are somewhat weaker and their frequency is higher. As those are more frequent, it might be easier to detect. Most remarkable are hidden jets that take place in some powerful supernovae [349]. Those jets choked inside supernovae and are difficult to observe directly. Their unique GW signature might enable us to confirm their existence. The energies of choked supernovae jets are comparable to those of long GRB jets [349]. As supernovae are much more frequent than GRBs it might be expected to detect these signals first [348]. In particular, the GWs from a powerful hidden jet in a supernovae are likely stronger than the GWs from the collapse event itself.
@@ 3.2.2 Populations and formation channels of GW sources
Main contributors: Elisa Bortolas, Manuel Arca Sedda, Pau Amaro Seoane, Stefano Benetti, Marica Branchesi, Adam Burrow, Enrico Cappellaro, Roberto Della Ceca, Ines Francesca Giudice, Elisabeth-Adelheid Keppler, Chiaki Kobayashi, Valeriya Korol, Francesco Longo, Ferdinando Patat, Sourav Roy Chowdhury, Roberto Serafinelli, Paola Severgnini, Martina Toscani, David Vartanyan
@@
Asymmetric NSs non axisymmetric rotating NS $1-4$?
@@
Stellar black-hole binaries – The detection of GWs by the LIGO-Virgo-Kagra (LVK) collaboration, led by the discovery of GW150914 [1], provided proof of BHs, enriching our knowledge of known merging BHs up to the present population of nearly 100 objects with masses above 30 M. This was well above the limit set by previous observations, supporting the evidence for the existence of binary BHs, and for the formation of new BHs as by-product of a BH mergers. Such a catalogue of events [55] allows placing constraints on the cosmic population of merging BHs. Among all observed mergers, there are several intriguing sources with properties that clearly deviate from the overall distributions. One of them is GW190521 [353, 354], the first GW source to leave behind an IMBH with a mass of $\sim 150M_{\odot}$.
@@
Finally, identifying the location of the merger would definitely help to pin down the formation scenario. Combining observations performed with different detectors can significantly help reduce the localisation area, possibly making it easier to assess the merger birth site (e.g. the properties of host galaxies and the location within the galaxies) at least on a statistical basis.
@@
White dwarf and neutron star binaries – Mergers of WDs and NSs are extremely important for understanding the origin of elements [445]. NS mergers are proposed for one of the astrophysical sites of the rapid neutron capture process [446], which has been confirmed by the detection of GW and EM transient ‘kilonova’ [447]. WDs mergers are thought to become sub-Chandrasekhar-mass Type Ia supernovae [448, 449], which are responsible for at least some of the Fe production particularly in dwarf galaxies [450]. Compact stellar systems composed of WDs and NSs are guaranteed GW sources, as some of them have been already discovered with electromagnetic (EM) telescopes (see also Section 3.3.1). Some of the shortest period and better characterised of the known binaries have been also proposed for testing space-based GW detectors to maximise their scientific output (e.g. [451, 452, 453, 454]). Some of such systems, known in literature as verification binaries, can be detected by LGWA [10], and offer the exciting opportunity to study the astrophysics of compact sources using both their GW and EM emission (e.g. [455, 456, 457, 458, 459]).
@@
From a theoretical perspective, compact binaries are expected to be several orders of magnitude more abundant than currently observed by EM telescopes (e.g. [460, 461]), with binaries composed of two WDs being the most abundant in the Milky Way and its immediate neighbourhood. Their formation involves up to several mass transfer episodes; at least one of these episodes has to be unstable leading to a common envelope, a phase during which the binary can shrink dramatically [462, 449, 463, 464, 465]. Even though this is one of the least understood phases in binary evolution, the fact that we do observe close double compact objects and mergers is a strong indication that something like a common envelope phase happens [466]. Based on these observations, several theoretical modelling efforts are being made (e.g. [467, 468]). Thus, the demographics of compact binaries observed via GW radiation will provide opportunities to learn new physics and answer key scientific questions related to the formation and evolutionary processes of close binary systems. Besides the common envelope physics, this includes questions related to the stability and efficiency of mass and angular momentum transfer, tides, and accretion onto compact objects, as well as details of their destruction in supernova explosions [for a review see 469, 470].
@@
At frequencies $<1\,$mHz, binary population synthesis simulations predict ${\cal O}(10^{7})$ double WDs, followed by ${\cal O}(10^{5})$ NSWDs in the Milky Way alone [469]. Studying these binary populations with EM facilities has proven to be technically challenging due to the compact size of these binaries and WD/NS stars themselves, and so GW detectors – characterised by different selection effects – are ideally suited for discovering and characterising them in bulk [471, 472]. Importantly, GW selection effects enable the study of EM-dim double compact objects at unprecedentedly large distances throughout the Milky Way, including its most massive satellites and possibly reaching out as far as the Andromeda galaxy, which otherwise impossible with EM telescopes [473, 474, 475, 476].
@@
Supernovae Type Ia as probe of white-dwarf binaries – SNe Ia are believed to originate from WDs which, due to mass transfer or generally, interactions in close binary systems, grow above the Chandrasekhar limit ($M_{Ch}\sim 1.4\,M_{\odot}$) and are incinerated by a thermonuclear explosion under degenerate matter conditions. This idea dates back to the early ’70 [477] but the actual nature of the secondary star remains undetermined. Two alternative scenarios have been proposed: the donor is a hydrogen or helium main sequence or a giant star (the single degenerate scenario, or SD), or another WD (the double degenerate scenario, or DD), which would lose orbital momentum via GW radiation and eventually merge. In the case of SD systems, mass transfer onto the accreting WD occurs when the secondary donor star fills its Roche lobe. This sets a lower limit for the separation between the two stars and the orbital period (of the order of 10 minutes for the most compact, non-degenerate donors [478]). In this configuration, the GW emission is weak and does not significantly affect the binary evolution. On the contrary, in DD systems the orbital radius naturally shrinks as an effect of angular momentum loss via GW emission [449]. Because of the small tidal radius of WDs, this allows a minimum orbital period of tens of seconds (see below). Since the intensity of the GW emission increases with increasing frequency, this is also expected to produce a much stronger signal, falling in the 0.01–0.1 Hz domain, right in the frequency range where LGWA’s sensitivity is at its best. This difference in orbital periods creates a clear-cut case, in which GW detection can act as an effective discriminant between the two channels. Using Kepler’s third law relating the period $P$ and the semi-major orbital axis $a$, and considering that the GW frequency is given by $f_{GW}=2/P$, it can be readily shown (see for instance [479]) that
@@
The SN Ia rate in the Galaxy is well known from direct observations in the local universe: $r_{\rm Ia}$=(5.4$\pm$1.2)$\times$10-3 yr-1 [483]. This corresponds to an event every $\sim$200 yr, which is definitely too low to provide a significant chance of detection of a WD binary merger during the lifetime of a GW campaign. However, in order to understand the progenitor of SN Ia, statistical approach can be followed. If the bulk of SNe Ia comes from DD systems, there needs to be enough such systems with the suitable properties (mass, merging time) to account for the observed SN Ia rate. Let us assume that $n(M_{\rm c},f_{\rm GW})$ is the probability density function describing the distribution of DD systems. Let us then also indicate with $M_{\rm c,l}$ and $M_{\rm c,h}$ the lower and upper chirp mass limits for having a total mass $M_{1}+M_{2}\gtrsim M_{\rm Ch}$. These limits are set considering the maximum and minimum mass of a C-O WD: $\sim 1.0M_{\odot}$ [484] and $\sim 0.3M_{\odot}$ [485] ($M_{1}=0.4M_{\odot}$, $M_{2}=1.0M_{\odot}$ and $M_{1}=M_{2}=1M_{\odot}$, corresponding to an approximate chirp mass range 0.5-0.9 $M_{\odot}$). Finally, let us indicate with $f_{\rm GW,l}$ the lower limit of the frequency range covered by the GW detector, which corresponds to the maximum merging time $t_{\rm m,h}$ and maximum orbital period $P_{h}$. With these settings, the total expected number of suitable DD systems found by the GW detector is:
@@
Since $r_{Ia}$ is well known (see above), the argument can be reversed and one can predict $N_{*}$. For illustration purposes, using this relation and Eq. (3.6), we have compiled Table 4 for some representative values of $P_{h}$. For the sake of simplicity, we have assumed $M_{1}=M_{2}=0.7M_{\odot}$ ($M_{c}\simeq 0.6M_{\odot}$). We note that the above calculations are a good approximation only for $P\gtrsim 100$ s. For lower periods (50-100 s, depending on the WD masses) Roche lobe overflow ensues. This leads to a merger timescale which is much shorter than what is predicted by Eq. (3.6).
@@
Considering the SNs observed by ZTF and ATLAS from 2018 to 2023, we estimate the number per year of SN Ia within 10, 20 and 100 Mpc by summing the SNe classified as Type Ia and assuming that, on the basis of percentage of Type Ia with respect to CCSNe, 40% of unclassified SNe are Type Ia. The numbers corrected for the effective sky areas observed by ATLAS and ZTF are reported in the last row of table 5.
@@
       10 Mpc        30 Mpc        100 Mpc
       Observed SN Ia        $1\pm 0.4$        $17.3\pm 1.7$        $86.7\pm 3.8$
       Observed Unclassified SN        $8.5\pm 1.2$        $31\pm 2.3$        $112\pm 4.3$
       Corrected SN Ia        $6.8\pm 1.4$        $46\pm 4.0$        $200\pm 8.5$
Table 5: The table shows the number of SNe Ia per year. The first row gives the number per year of SN Ia observed mainly by the ZTF and ATLAS surveys considering the observations from 2018 to 2023. The second row gives the rate of unclassified SNe coming from the same surveys. The last row gives the expected rate of SNe Ia observed by an all-sky monitor. This rate is obtained by adding the number of observed SN Ia (first row) to 40% of unclassified SNe (second row), where 40% is the percentage of SNe Ia among classified SN. This value is then corrected by the surveyed effective areas of about 65%.
Considering a merging frequency of 0.1 Hz for a 1+1 $M_{\odot}$ DWD system and the LGWA DWD horizon distance that continuously increases with observation time up to 10 yr, we estimate a total number of about 160 DWD mergers possibly detectable by LGWA over the 10 years of mission lifetime. This number assumes a conservative 50% percent duty cycle for optical surveys (due to observation only during darkness). It does not take into account that some SN Ia are not detected in the field of view of ZTF/ATLAS due to obstruction of the EM radiation.
This detection rate enables the interesting possibility of measuring DWD systems in galaxies that in the past hosted a SN Ia. In that case, host galaxy distances derived from the GW signal might be used as an independent probe to calibrate the SN distance for cosmology.

### v2

@@ 3.2.1 Astrophysical explosions and matter effects on waveforms
In the last two decades, significant efforts have been made in investigating the GW emission from CCSNe (see [316, 317, 318, 319, 320] for recent reviews). Non-spherical explosions are required to produce GWs, which happens during three distinct phases: a) immediately at bounce (through time-dependent rotational flattening and post-shock convection, lasting tens of ms); b) the growing neutrino-driven turbulence (lasting 50–200 ms); and through c) PNS modal oscillations [321, 322, 323, 324, 325, 326, 327, 328, 316]. These phases, characterised by GW emission in the $\sim$100–1000 Hz frequency range, are accompanied by an additional low-frequency ($\lesssim$50 Hz) signal, henceforth known as the GW linear memory [329, 330] (see also Sec. 3.2.2) that is generated through aspherical, large-scale matter ejecta motions coupled with anisotropic neutrino emission [331, 332, 333, 334, 335, 336]. The neutrino memory signal shows a more secular time-evolution than the matter component, which is due to the cumulative time-integral of the anisotropy-weighted neutrino luminosity [337]. Sustained turbulent accretion in more massive progenitors results in higher neutrino luminosities and generally more anisotropic explosions. As with the matter component, the neutrino component is more pronounced for delayed explosions of models with higher compactness [338].
@@
Large-scale SN ejecta asymmetries (where jet-like structures may emerge) are expected (see e.g. [339, 340, 341, 342, 343]) and are seen observationally using spectropolarimetry [344, 345, 346, 347, 348, 349], but their GW emissions are associated with frequencies that are beyond the LGWA observation band (100–1000 Hz) [350, 351, 352, 353] (see also Fig. 28).
@@
Figure 28: Gravitational-wave spectra $2\sqrt{f}|\tilde{h}(f)|$ for CCSNe with different stellar masses as a function of time projected against strain noise curves $\sqrt{S_{n}}$ for various GW detectors (adapted from [319]). The simulations, extending out to $\sim$6 seconds after the bounce, are unable to constrain the spectrogram in the region 0.1–1.0 Hz. The sources are assumed to be at a distance of 10 kpc.
@@
While recent simulations do not cover the decihertz band (see for instance [319] and Fig. 28 for 10 kpc event signals) the linear and nonlinear memory effects can lead to detectable GW signals below 1 Hz [354, 335]. In this way, even with LGWA, one would be able to study the GW emission generated by the neutrino processes [355, 334, 319]. As discussed above, the range $>$50 Hz primarily arises from matter asymmetries while the range $<$10 Hz carries information regarding the anisotropy of the neutrino emission. Although these frequency ranges may overlap in reality, they are, in principle, distinguishable. The polarization of the signal provides insights into the asymmetries of the event along with the orientation of the source with respect to the line of sight. Notably, a simultaneous detection of the low-frequency memory signal by LGWA and the high-frequency component by terrestrial detectors would offer an invaluable, unprecedented understanding of CCSNe. The memory signal would provide information about the progenitor star while frequencies above 100 Hz would provide insight into the newly formed PNS. Although the exploration of lower-frequency CCSNe signals and their detectability has only begun in recent years, significant progress has been made in extending simulations to several seconds post-bounce. The near-future goal of extending these simulations to tens of seconds holds promise for producing consistent results at low frequencies.
@@
The acceleration of a relativistic jet is another source of gravitational radiation [356, 357, 358]. The resulting GW signal is a memory-type signal, namely, they take the form of a step leaving a permanent deformation of spacetime after they pass. The detailed features of the signal depend on the acceleration (and possible deceleration) of the jet, on its overall duration and orientation relative to the observer [359, 360, 361]. The overall amplitude of the GW signal is, as expected on dimensional grounds, of order $GE_{\rm jet}/c^{4}d$, where $E_{\rm jet}$ is the energy of the jet and $d$ the distance of the source. The frequency is the smaller of the inverse acceleration time and the jet duration. Natural candidates for jet-GWs are long GRB jets whose typical energy is of order $10^{51}$ erg and typical frequency of order tens of deciHz. Short GRB jets are somewhat weaker and have higher frequency. As those are more frequent, it might be easier to detect. Most remarkable are hidden jets that take place in some powerful supernovae [362]. Those jets are choked inside supernovae and are difficult to observe directly. Their unique GW signature might enable us to confirm their existence. The energies of choked supernovae jets are comparable to those of long GRB jets [362]. As supernovae are much more frequent than GRBs it might be expected to detect these signals first [361]. In particular, the GWs from a powerful hidden jet in supernovae are likely stronger than the GWs from the collapse event itself. A Jet-GW signal from a galactic supernova will have a strain of $\gtrsim 10^{-21}$ in the frequency range of LGWA.
@@ 3.2.2 Populations and formation channels of GW sources
Main contributors: Elisa Bortolas, Manuel Arca Sedda, Pau Amaro Seoane, Stefano Benetti, Marica Branchesi, Adam Burrows, Enrico Cappellaro, Roberto Della Ceca, Kiranjyot Gill, Ines Francesca Giudice, Elisabeth-Adelheid Keppler, Chiaki Kobayashi, Valeriya Korol, Francesco Longo, Ferdinando Patat, Sourav Roy Chowdhury, Roberto Serafinelli, Paola Severgnini, Martina Toscani, David Vartanyan
@@
Asymmetric NSs Non-axisymmetric rotating NS $1-4$?
@@
Stellar black-hole binaries – The detection of GWs by the LIGO-Virgo-Kagra (LVK) collaboration, led by the discovery of GW150914 [1], has provided proof of BHs, enriching our knowledge of known merging BHs up to the present population of nearly 100 objects with masses above 30 M. This was well above the limit set by previous observations, supporting the evidence for the existence of binary BHs, and for the formation of new BHs as by-product of a BH mergers. Such a growing catalogue of events [54] allows constraints to be placed on the cosmic population of merging BHs. Among all observed mergers, there are several intriguing sources with properties that clearly deviate from the overall distributions. One of them is GW190521 [366, 367], the first GW source to leave behind an IMBH with a mass of $\sim 150M_{\odot}$.
@@
Finally, identifying the location of the merger would definitely help to pin down the formation scenario. Combining observations performed with different detectors can significantly help reduce the localization area down to a few arcmin2 (see section 2.2.2), possibly providing clues about the merger birth site, at least on a statistical basis and for relatively closeby sources.
@@
White-dwarf and neutron-star binaries – Mergers of WDs and NSs are extremely important for understanding the origin of elements [458]. NS mergers are proposed for one of the astrophysical sites of the rapid neutron capture process [459], which has been confirmed by the detection of GW and EM transient ‘kilonova’ [460]. WDs mergers are thought to become sub-Chandrasekhar-mass Type Ia supernovae [461, 462], which are responsible for at least some of the Fe production particularly in dwarf galaxies [463]. Compact stellar systems composed of WDs and NSs are guaranteed GW sources, as some of them have been already discovered with electromagnetic (EM) telescopes (see also Section 3.3.1). Some of the shortest period and better characterised of the known binaries have also been proposed for testing space-based GW detectors to maximise their scientific output (e.g. [464, 465, 466, 467]). Some of such systems, known in literature as verification binaries [e.g., 464], can be detected by LGWA [10]. They are guaranteed multi-messenger sources offering an opportunity to study the astrophysics of compact sources using both their GW and EM emission (e.g. [468, 469, 302, 470, 471]). They would also be ideal calibration sources for LGWA given their known location and well-modeled GW amplitudes eventually measured precisely by LISA (see section 2.1.4.
@@
From a theoretical perspective, compact binaries are expected to be several orders of magnitude more abundant than currently observed by EM telescopes (e.g. [472, 473]), with binaries composed of two WDs being the most abundant in the Milky Way and its immediate neighbourhood. Their formation involves up to several mass transfer episodes; at least one of these episodes has to be unstable leading to a common envelope, a phase during which the binary can shrink dramatically [474, 462, 475, 476, 477]. Even though this is one of the least understood phases in binary evolution, the fact that we do observe close double compact objects and mergers is a strong indication that something like a common envelope phase happens [478]. Based on these observations, several theoretical modelling efforts are being made (e.g. [479, 480]). Thus, the demographics of compact binaries observed via GW radiation will provide opportunities to learn new physics and answer key scientific questions related to the formation and evolutionary processes of close binary systems. Besides the common envelope physics, this includes questions related to the stability and efficiency of mass and angular momentum transfer, tides, and accretion onto compact objects, as well as details of their destruction in supernova explosions [for a review see 481, 482].
@@
At frequencies $<0.1\,$mHz, binary population synthesis simulations predict ${\cal O}(10^{7})$ double WDs, followed by ${\cal O}(10^{5})$ NSWDs in the Milky Way alone [for a review see 481]. Studying these binary populations with EM facilities has proven to be technically challenging due to the compact size of these binaries and WD/NS stars themselves, and so GW detectors – characterised by different selection effects – are ideally suited for discovering and characterising them in bulk [483, 484]. Importantly, GW selection effects enable the study of EM-dim double compact objects at unprecedentedly large distances throughout the Milky Way, including its most massive satellites and possibly reaching out as far as the Andromeda galaxy, which otherwise impossible with EM telescopes [485, 486, 487, 488]. Recent binary population synthesis studies forecast between 2 and 200 double WDs in the Milky Way with orbital periods of less than 200 seconds (GW frequencies greater than $10^{-2}$ Hz) that LGWA could potentially capture in the inspiral phase [469, 489, 490, 491, 492, 493]. Observation-based forecasts are even more optimistic, predicting up to 1000 inspiralling double WDs in the same frequency regime in the Milky Way alone [494, 255, 484].
@@
Supernovae Type Ia and white-dwarf binaries – SNe Ia are believed to originate from WDs which, due to mass transfer or generally, interactions in close binary systems, grow above the Chandrasekhar limit ($M_{Ch}\sim 1.4\,M_{\odot}$) and are incinerated by a thermonuclear explosion under degenerate matter conditions. This idea dates back to the early ’70 [495] but the actual nature of the secondary star remains undetermined. Two alternative scenarios have been proposed: the donor is a hydrogen or helium main sequence or a giant star (the single degenerate scenario, or SD), or another WD (the double degenerate scenario, or DD), which would lose orbital momentum via GW radiation and eventually merge. In the case of SD systems, mass transfer onto the accreting WD occurs when the secondary donor star fills its Roche lobe. This sets a lower limit for the separation between the two stars and the orbital period (of the order of 10 minutes for the most compact, non-degenerate donors [496]). In this configuration, the GW emission is weak and does not significantly affect the binary evolution. On the contrary, in DD systems the orbital radius naturally shrinks as an effect of angular momentum loss via GW emission [462]. Because of the small tidal radius of WDs, this allows a minimum orbital period of tens of seconds (see below). Since the intensity of the GW emission increases with increasing frequency, this is also expected to produce a much stronger signal, falling in the 0.01–0.1 Hz domain, right in the frequency range where LGWA’s sensitivity is at its best. This difference in orbital periods creates a clear-cut case, in which GW detection can act as an effective discriminant between the two channels. Using Kepler’s third law relating the period $P$ and the semi-major orbital axis $a$, and considering that the GW frequency is given by $f_{GW}=2/P$, it can be readily shown (see for instance [497]) that
@@
The SN Ia rate in the Galaxy is well known from direct observations in the local universe: $r_{\rm Ia}$=(5.4$\pm$1.2)$\times$10-3 yr-1 [501]. This corresponds to an event every $\sim$200 yr, which is definitely too low to provide a significant chance of detection of a WD binary merger during the lifetime of a GW campaign.
However, in order to understand the progenitor of SN Ia, statistical approach can be followed. If the bulk of SNe Ia comes from DD systems, there needs to be enough such systems with the suitable properties (mass, merging time) to account for the observed SN Ia rate. Let us assume that $n(M_{\rm c},f_{\rm GW})$ is the probability density function describing the distribution of DD systems. Let us then also indicate with $M_{\rm c,l}$ and $M_{\rm c,h}$ the lower and upper chirp mass limits for having a total mass $M_{1}+M_{2}\gtrsim M_{\rm Ch}$. These limits are set considering the maximum and minimum mass of a C-O WD: $\sim 1.0M_{\odot}$ [502] and $\sim 0.3M_{\odot}$ [503] ($M_{1}=0.4M_{\odot}$, $M_{2}=1.0M_{\odot}$ and $M_{1}=M_{2}=1M_{\odot}$, corresponding to an approximate chirp mass range 0.5-0.9 $M_{\odot}$). Finally, let us indicate with $f_{\rm GW,l}$ the lower limit of the frequency range covered by the GW detector, which corresponds to the maximum merging time $t_{\rm m,h}$ and maximum orbital period $P_{h}$. With these settings, the total expected number of suitable DD systems found by the GW detector is:
@@
Since $r_{Ia}$ is well known (see above), the argument can be reversed and one can predict $N_{*}$. For illustration purposes, using this relation and Eq. (3.6), we have compiled Table 4 for some representative values of $P_{h}$. For the sake of simplicity, we have assumed $M_{1}=M_{2}=0.7M_{\odot}$ ($M_{c}\simeq 0.6M_{\odot}$). We note that the above calculations are a good approximation only for $P\gtrsim 100$ s. For lower periods (50-100 s, depending on the WD masses) Roche lobe overflow ensues. Recent 3D hydrodynamical simulations show that for sufficiently massive primaries, the onset of unstable dynamical accretion of helium from the secondary white dwarf, just prior to the binary merger, heats up the helium shell on the primary white dwarf. As a result, a thermonuclear runaway occurs, igniting a helium detonation on the surface of the primary white dwarf and eventually leading to the ignition of the carbon-oxygen core [e.g., 504, 505, 506]. This would lead to a merger timescale that is much shorter than what is predicted by Eq. (3.6).
@@
[…]
The numbers thus estimated are conservative. A non-negligible fraction of SNe are not classified and could potentially increase the SN Ia number, and detection incompleteness can start to arise at large distance due to the limited surveys’ sensitivity. However, both biases can be considered negligible within 20 Mpc.
@@
      10 Mpc       20 Mpc       50 Mpc       100 Mpc
      Observed rate       $0.7\pm 0.3$       $1.8\pm 0.5$       $17.3\pm 1.7$       $86.7\pm 3.8$
      Corrected rate       $1\pm 0.4$       $2.7\pm 0.7$       $26\pm 2.6$       $130\pm 5.7$
Table 5: The table shows the number of SNe Ia per year. The first row gives the number per year of SN Ia based on the observations of ZTF and ATLAS. The second row gives the rate corrected by the effective sky areas of about 65% surveyed by ZTF and ATLAS.
Considering a merging frequency of 0.1 Hz for a 1+1 $M_{\odot}$ DWD system and the LGWA DWD horizon distance that continuously increases with observation time up to 10 yr, we estimate a total number of about 20 DWD mergers producing SNe Ia possibly detectable by LGWA over the 10 years of mission lifetime. This number does not take into account that some SN Ia are not detected by ZTF/ATLAS due to obstruction of the EM radiation, but such systems would not escape detection in GWs.
These detections enable the interesting possibility of measuring DWD systems in galaxies that in the past hosted a SN Ia. In that case, host galaxy distances derived from the GW signal might be used as an independent probe to calibrate the SN distances for cosmology. It should also be noted that DWD mergers cannot produce a SN Ia when the total mass of the binary lies below the Chandrasekhar limit. This means that many more DWD mergers should be expected to be observable with LGWA than the number inferred from SN Ia observations, and those could be used equally well to calibrate SN distances.